Problem 5-46 EAR of Add-On Interest Loan (LG5-7, LG5-8) 


To borrow $1,350, you are offered an add on interest loan at 8.7 percent with 12 monthly payments. Compute the 12 equal payments. (Round your answer to 2 decimal places.) 


Use the amount you borrowed and the monthly payments you computed to calculate the APR of the loan. Then, use that APR to compute the EAR of the loan. (Do not round intermediate calculations and round your answer to 2 decimal places.)  
Answer #1

Answer (a):

Equal Payments $ 122.29

Working:

Add-on interest loan:

Amount borrowed = $1350

Interest amount = 1350 * 8.7% = $117.45

Monthly payments = (1350 + 117.45) / 12 = $122.2875 or $122.29

Answer (b):

EAR 16.87 %

Working:

Monthly payment as calculated above = PMT = $122.2875

PV = $1350

NPER in months = 12

Monthly interest = RATE(nper, pmt, pv, fv, type) = RATE(12,122.2875,-1350, 0, 0) = 1.3073429439459%

APR = 1.3073429439459% * 12 = 15.688115327351%

EAR = (1 + 15.688115327350%/12)^12 - 1

= 16.87%